OpenStudy (anonymous):

Using the equation y equals x divided by five, what is the value of y when x is twenty? y = x over 5 x y 5 1 15 3 20 ? 40 8 4 5 6 7

OpenStudy (hoblos):

y=x/5 for x=20, y= 20/5= ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

thx